OpenStudy (anonymous):

Simplifying this expression? 5√24 * 3√10?

OpenStudy (snowfire):

You can combine the insides of the square roots since they are being multiplied: 5sqrt(24)*3sqrt(10)=15sqrt(240)

OpenStudy (anonymous):

can you simplify it further or is 240 it..? Cause thats where i got confused. I got 15√240 as well.

OpenStudy (snowfire):

Then you can separate the 240 into its factors: 240=2*120=2*2*60=2*2*2*30=2*2*2*2*15=2*2*2*2*3*5=2^4*15 Since this is all within a square root, every pair can be brought out: sqrt(2^4)=sqrt((2^2)*(2^2))=4 And you leave anything inside that has no pairs.

OpenStudy (anonymous):

okay thankyou

OpenStudy (snowfire):

Yep no problem